# Non-Iterative Symbolic-Aided Chain-of-Thought for Logical Reasoning

### v1

@@ Abstract
This work introduces Symbolic-Aided Chain-of-Thought (CoT), an improved approach to standard CoT, for logical reasoning in large language models (LLMs). The key idea is to integrate lightweight symbolic representations into few-shot prompts, structuring the inference steps with a consistent strategy to make reasoning patterns more explicit within a non-iterative reasoning process. By incorporating these symbolic structures, our method preserves the generalizability of standard prompting techniques while enhancing the transparency, interpretability, and analyzability of LLM logical reasoning. Extensive experiments on four well-known logical reasoning benchmarks—ProofWriter, FOLIO, ProntoQA, and LogicalDeduction, which cover diverse reasoning scenarios—demonstrate the effectiveness of the proposed approach, particularly in complex reasoning tasks that require navigating multiple constraints or rules. Notably, Symbolic-Aided CoT consistently improves LLMs’ reasoning capabilities across various model sizes and significantly outperforms conventional CoT on three out of four datasets, ProofWriter, ProntoQA, and LogicalDeduction.
Non-Iterative Symbolic-Aided Chain-of-Thought
@@
Phuong Minh Nguyen and Tien Huu Dang and Naoya Inoue RebelsNLU Lab, IS school Japan Advanced Institute of Science and Technology {phuongnm,s2310417,naoya-i}@jaist.ac.jp Correspondence: phuongnm@jaist.ac.jp
@@ 1 Introduction
Notably, as model size increases, prompting techniques that incorporate intermediate reasoning steps consistently outperform standard input–output prompting methods. This reasoning-based prompting strategy is known as Chain-of-Thought prompting (CoT; Wei et al. (2022)). For in-context learning, the CoT reasoning technique has demonstrated compelling results across a variety of NLP tasks Wei et al. (2022); Zhou et al. (2023). Despite recent advancements, applying LLMs to logical reasoning tasks still faces several critical challenges, including conflicts between pretrained knowledge and counterfactual assumptions (Ortu et al., 2024), failures on cyclic inference graphs (Zhang et al., 2025), and planning errors during the solving process Ye et al. (2023).
@@
In this study, we focus on the second approach, which aims to improve LLM logical reasoning without relying on any external symbolic reasoner or programming language. Building on insights from recent works Sun et al. (2024); Qi et al. (2025); Zhang et al. (2025), we target the challenge of designing a mechanism to systematically decompose complex logical reasoning tasks into simpler subtasks that can be solved by the inherent understanding ability of LLMs in a single inference pass. Closely related to our work, Xu et al. (2024a); Feng et al. (2024) recently introduced LLM-based frameworks that leverage first-order logic—a strict formal language with well-defined syntax and semantics—to support faithful logical reasoning. However, these prior studies Sun et al. (2024); Xu et al. (2024b); Zhang et al. (2025) primarily rely on self-refinement or iterative (multi-turn) generation, where each turn solves a sub-task and its output feeds into the next. They overlook non-iterative reasoning. To bridge this gap, we investigate non-iterative (single-turn) reasoning generation, allowing a more direct and fair comparison with CoT prompting.
@@
We introduce Symbolic-Aided CoT, a novel variant of CoT Wei et al. (2022), designed to leverage symbolic representations to enhance the non-iterative logical inference capabilities of LLMs (Figure 1). In conventional CoT, intermediate reasoning steps are provided in few-shot demonstrations as unstructured text, enabling LLMs to approximate the logical reasoning process. However, relying solely on textual descriptions for complex reasoning introduces ambiguity, as the inherent vagueness of natural language limits LLMs’ ability to generalize precise reasoning steps. Our core idea is to integrate lightweight symbolic structures into few-shot prompts, making the inference steps more transparent and structured, while simultaneously strengthening the induction-head behavior of LLMs Olsson et al. (2022). Specifically, our Symbolic-Aided CoT prompting framework explicitly outlines essential reasoning substeps: rule matching—selecting constraint rules that align with the current state of inference, new premise inference—applying the selected rules to generate new premises, and knowledge base (KB) updating—appending the newly inferred premises to the KB set. By incorporating these symbolic structures, our method preserves the flexibility and general applicability of standard prompting techniques while enhancing both the interpretability and analyzability of LLM reasoning behavior. Empirical evaluations across four reasoning QA benchmarks, ProofWriter, FOLIO, ProntoQA, and LogicalDeduction, demonstrate the effectiveness of our approach, particularly in scenarios involving complex reasoning that requires navigating multiple rules and constraints. Remarkably, our Symbolic-Aided CoT, when applied with open-source LLMs (e.g., Qwen3), achieves performance comparable to that of a complex multi-turn reasoning framework Sun et al. (2024) built on the powerful GPT-4 model, particularly on the ProofWriter, ProntoQA, and LogicalDeduction datasets.
@@ 2.1 Background
In the second approach, the original logical reasoning problem is decomposed into a sequence of subtasks. Each subtask is solved individually, and the process iterates over multiple turns until a specified stopping condition is satisfied, ultimately producing the final result: $\mathrm{Loop([\mathrm{LLM}^{\textrm{subtask}_{t}}(\mathcal{R},Q)]_{t})}$ where $\textrm{subtask}_{t}$ may represent an arbitrary LLM-based unit function such as rule matching, rule inference, or new premise scoring, among others. These sub-tasks are typically carefully designed and arranged with sequential dependencies, with the aim of mitigating hallucinations.
@@ 2.2 Related Works
In this line of work, the main challenge lies in transforming the description of constraints into logical formulas, $\mathrm{LLM}^{\textrm{lang2logic}}$, effectively functioning as a semantic parser. More specifically, Yao et al. (2023) introduced a method to enrich CoT prompting by leveraging results from external APIs, invoking functions to retrieve supplementary information. Building on this idea, Gao et al. (2023) proposed an approach that augments the intermediate reasoning steps of CoT reasoning through a runtime environment (e.g., a Python interpreter), which has proven particularly effective for mathematical reasoning tasks. Further, Pan et al. (2023); Ye et al. (2023); Olausson et al. (2023) enhanced the performance of logical reasoning by translating all constraints in a sample into logical forms, completing a logical program, and solving it using an independent symbolic reasoner. In addition, Xu et al. (2024a) strengthened the logical-form parsing process of the open-source LLMs by leveraging fine-tuning on a large-scale, curated dataset. In contrast, our method does not rely on any external symbolic solver; instead, it integrates symbolic syntax directly into the reasoning process, aiming to enable the LLM itself to reason as a symbolic solver.
@@
This approach leverages LLMs directly by designing frameworks that decompose complex reasoning tasks into smaller, manageable subtasks—such as rule selection, premise derivation, and scoring—to enhance overall reasoning robustness (Zhang et al., 2025; Feng et al., 2024; Sun et al., 2024). In particular, Sun et al. (2024); Xu et al. (2024a) introduced a framework that enables LLMs to uncover hidden premises and integrates scoring components through a multi-turn reasoning process. Similarly, Zhang et al. (2025) proposed a framework that incrementally refines reasoning via three subcomponents: Proposer, Verifier, and Reporter. Finally, Feng et al. (2024) presented LoGiPT, a method that enhances LLMs’ ability to function as logical solvers by learning the reasoning process step by step through additional training on large-scale data collected from the reasoning traces of external symbolic solvers. Compared to our work, LoGiPT similarly performs step-by-step reasoning as a logical solver and can produce a proof tree at the end; however, our method achieves competitive performance without the need for any additional training.
@@ 3 Methodology
In this section, we present our Symbolic-Aided CoT method, its variants, and the motivation behind it in comparison with baseline prompting techniques: Standard - which directly provides an answer without any reasoning - and CoT (Wei et al., 2022) - which produces an answer accompanied by step-by-step reasoning. All of these prompting methods are augmented with a hard-selected few-shot examples included in the prompt Qi et al. (2025). All prompts are directly fed forward through the LLM to obtain the final predicted answer in a single turn:
$A^{out}=\mathrm{LLMs}(\mathrm{prompting}(\mathcal{R},Q))$ (1)
In this setup, the LLM is solely responsible for generating the desired answer given the contextual input. The prompting component consists of a few-shot template designed to help the LLM understand the task description while leveraging its own knowledge to reason over the list of provided rules in the contextual information. For clarity, the templates for Standard and CoT prompting are shown in the first two rows of Table 1.
@@ 3.1 Symbolic-Aided CoT
Inspired by recent works Sun et al. (2024); Zhang et al. (2025), we formulate logical reasoning tasks into three fundamental sub-tasks (reasoning operators): rule matching, rule inference, and knowledge base updating. Previous frameworks, such as DetermLR Sun et al. (2024) and CR Zhang et al. (2025), were also built on carefully designed unit operators, integrating them with procedural programming to process the outputs of LLMs. The key difference between our Symbolic-Aided CoT and these approaches is that Symbolic-Aided CoT is conceived entirely as an LLM-driven program. In our design, the LLM is expected to learn the flow of the logical reasoning process from a few-shot demonstration. To this end, the LLM has full visibility of all sub-reasoning steps and autonomously decides which step to execute next. The overview of our Symbolic-Aided CoT is presented in the third row of Table 1, which illustrates the instruction text, list of rules, question, and reasoning-flow examples. For a clearer explanation, we elaborate on the two gray blocks shown in this table, which pertain to rule tagging and reasoning operators in our Symbolic-Aided CoT.
@@
This demonstration serves as the core example that enables LLMs to learn, in context, the pattern for solving logical reasoning tasks. We use a set of symbols, similar to those in programming languages, to represent the inference flow (see Figure 2). At each reasoning step, the LLM selects the relevant rules and premises from the current knowledge base (KB) to infer new premises. Each newly inferred premise is then appended to the KB for use in the next inference step. A breadth-first search strategy is applied to traverse the nodes (premises), as illustrated in Figure 2. Inspired by how humans solve such tasks, we maintain a KB state to prevent cyclical inference loops: if a newly inferred premise already exists in the KB, it is not added again. All patterns of rule selection, inference, and KB updating are implicitly conveyed within the demonstrations provided in the few-shot prompts, allowing the LLM to internalize these reasoning steps.
Figure 2: Examples of reasoning flows in the Symbolic-Aided CoT demonstrations.
@@ 4.1 Experimental Setup
We conducted our experiments primarily on open-source LLMs, including Llama-3.1-8B-Instruct Grattafiori et al. (2024) and the Qwen-3 models Yang et al. (2025). Specifically, we aim to evaluate the effectiveness of our Symbolic-Aided CoT compared to standard CoT and standard prompting on these LLMs. In addition, we performed extensive experiments on the Qwen-3 models across various sizes - 1.7B, 4B, 8B, and 32B - to further assess the scalability and effectiveness of our proposed prompting method. Finally, for comparison with previous approaches, we also conducted extensive experiments on a powerful closed-source LLM, GPT-4 Achiam et al. (2023), to evaluate our proposed method. For open-source models, we use greedy decoding to generate the answer (i.e., the token with the maximum logit is picked).
@@ 4.2 Main Results
These results demonstrate the clear superiority of our proposed method over basic CoT Wei et al. (2022) on three datasets, ProofWriter, ProntoQA, and LogicalDeduction. Notably, on the ProofWriter dataset, Symbolic-Aided CoT significantly outperforms CoT, achieving improvements of $15\%$ and $1\%, 22$$2\%$ on the Qwen-3-8B, Qwen-3-14B, and Llama3.1-8B-Instruct models, respectively. In addition, the improvement is clearly observed on the LogicalDeduction and ProntoQA datasets across three open-source LLMs. These findings further highlight that the degree of improvement varies across different LLMs. The effectiveness of Symbolic-Aided CoT largely depends on each model’s ability to understand logical relationships and recognize logical matching patterns embedded within the few-shot prompts. Moreover, our method is simple yet effective, achieving competitive performance compared to prior works such as Logic-LM Pan et al. (2023) and DetermLR Sun et al. (2024), even when those methods are supported by more powerful LLMs like GPT-4.
@@
Finally, we evaluate our proposed method in the setting that uses GPT-4 as the backbone LLM for the reasoning task (last row of Table 2). Compared to the SymbCoT framework Xu et al. (2024a), our method achieves superior performance on ProntoQA but lower performance on other datasets. This difference can be attributed to SymbCoT’s use of complex interactive reasoning sub-steps—such as translator, planner, solver, and verifier—each supported by carefully designed prompts tailored to the specific sub-step and logical reasoning task. Overall, our method surpasses the performance of previous methods on the ProntoQA dataset and achieves remarkable results on the ProofWriter, FOLIO, and LogicalDeduction datasets. These results demonstrate the robustness of our approach, even in the stringent setting that uses only non-iterative inference without the support of an external solver or multi-step inferences, such as the DetermLR Sun et al. (2024) or LogicLM Pan et al. (2023) approaches.
@@
Fine-tuned ID Qi et al. (2025) Supervised FT Non-Iterative $71.67$ $70.00$ - -
Standard ICL few-shot Non-Iterative $36.83$ $53.92$ $51.80$ $40.67$
CoT ICL few-shot Non-Iterative $44.83$ $\mathbf{56.86}$ $74.00$ $58.00$
Symbolic-Aided CoT (ours) ICL few-shot Non-Iterative $\mathbf{68.67}$ $55.88$ $\mathbf{89.00}$ $\mathbf{59.33}$
@@
Standard ICL few-shot Non-Iterative $60.00$ $62.25$ $80.80$ $57.00$
CoT ICL few-shot Non-Iterative $57.83$ $\mathbf{66.67}$ $95.80$ $72.67$
Symbolic-Aided CoT (ours) ICL few-shot Non-Iterative $\mathbf{78.67}$ $65.69$ $\mathbf{97.20}$ $\mathbf{77.33}$
@@
Standard ICL few-shot Non-Iterative $46.50$ $67.16$ $77.80$ $61.67$
CoT ICL few-shot Non-Iterative $62.67$ $\mathbf{74.02}$ $97.20$ $81.67$
Symbolic-Aided CoT (ours) ICL few-shot Non-Iterative $\mathbf{77.00}$ $65.20$ $\mathbf{97.80}$ $\mathbf{86.33}$
@@
CoT Sun et al. (2024) ICL few-shot Non-Iterative $67.41$ $67.65$ $91.00$ $73.33$
Logic-LM Pan et al. (2023) ICL few-shot Iterative +E.Solver $79.66$ $78.92$ $83.20$ $87.63$
DetermLR Sun et al. (2024) ICL few-shot Iterative +Programming $79.17$ $75.49$ $98.60$ $85.00$
SymbCoT Xu et al. (2024a) ICL few-shot Iterative +Programming $\mathbf{82.50}$ $\boldsymbol{83.33}$ $99.60$ $\mathbf{93.00}$
Symbolic-Aided CoT (ours) ICL few-shot Non-Iterative $77.09$ $74.51$ $\boldsymbol{100.00}$ $86.33$
@@
Furthermore, KB-tracking intuitively provides additional features to the hidden representation of premises, allowing the model to distinguish inferred premises from conditional premises in the constraint rules. In another aspect, the symbolic $\mathrm{Validate}$ function in our SymbolicCoT helps LLMs refer back to the original question in the context to select the appropriate premise for logical matching and producing the final answer.

### v2

# Non-Interactive Symbolic-Aided Chain-of-Thought for Logical Reasoning

@@ Abstract
This work introduces Symbolic-Aided Chain-of-Thought (CoT), an improved approach to standard CoT, for logical reasoning in large language models (LLMs). The key idea is to integrate lightweight symbolic representations into few-shot prompts, structuring the inference steps with a consistent strategy to make reasoning patterns more explicit within a non-interactive reasoning process. By incorporating these symbolic structures, Symbolic-Aided CoT preserves the generalizability of standard prompting techniques while enhancing the transparency, interpretability, and analyzability of LLM logical reasoning. Extensive experiments on four well-known logical reasoning benchmarks—ProofWriter, FOLIO, ProntoQA, and LogicalDeduction, which cover diverse reasoning tasks and scenarios—demonstrate the effectiveness of the proposed approach, particularly in complex reasoning tasks that require navigating multiple constraints or rules. Notably, Symbolic-Aided CoT consistently improves LLMs’ reasoning capabilities across various model sizes and significantly outperforms conventional CoT on three out of four datasets, ProofWriter, ProntoQA, and LogicalDeduction.
Non-Interactive Symbolic-Aided Chain-of-Thought
@@
Phuong Minh Nguyen and Tien Huu Dang and Naoya Inoue Japan Advanced Institute of Science and Technology {phuongnm,tiendh,naoya-i}@jaist.ac.jp Correspondence: phuongnm@jaist.ac.jp
@@ 1 Introduction
Notably, as model size grows, prompting methods that leverage intermediate reasoning steps consistently surpass standard input–output prompting methods. This reasoning strategy, known as Chain-of-Thought prompting (CoT; Wei et al. (2022)), relies on explicitly modeling the reasoning process. For in-context learning, the CoT reasoning technique has demonstrated compelling results across a variety of NLP tasks Wei et al. (2022); Zhou et al. (2023). Despite recent advancements, applying LLMs to logical reasoning tasks still faces several critical challenges, including conflicts between pretrained knowledge and counterfactual assumptions (Ortu et al., 2024), failures on cyclic inference graphs (Zhang et al., 2025), and planning errors during the solving process Ye et al. (2023).
@@
In this study, we focus on the second approach, which aims to improve LLM logical reasoning without relying on any external symbolic reasoner or programming language. Building on insights from recent works Sun et al. (2024); Qi et al. (2025); Zhang et al. (2025), we target the challenge of designing a mechanism to systematically decompose complex logical reasoning tasks into simpler subtasks that can be solved by the inherent understanding ability of LLMs in a single inference pass. Closely related to our work, Xu et al. (2024a); Feng et al. (2024) recently introduced LLM-based frameworks that leverage first-order logic—a strict formal language with well-defined syntax and semantics—to support faithful logical reasoning. However, previous studies Sun et al. (2024); Xu et al. (2024b); Zhang et al. (2025) primarily rely on self-refinement or interactive (multi-turn) generation, where each turn solves a sub-task and its output is passed to the next. They overlook non-interactive reasoning, in which the LLM performs the entire reasoning process without any assistance from external modules or sub-processes. To address this gap, we explore non-interactive (single-turn) reasoning generation, allowing a more direct and fair comparison with CoT prompting.
@@
We introduce Symbolic-Aided CoT, a novel variant of CoT Wei et al. (2022), designed to leverage symbolic representations to enhance the non-interactive logical inference capabilities of LLMs (Figure 1). In conventional CoT, intermediate reasoning steps are provided in few-shot demonstrations as unstructured text, enabling LLMs to approximate the logical reasoning process. However, relying solely on textual descriptions for complex reasoning introduces ambiguity, as the inherent vagueness of natural language limits LLMs’ ability to generalize precise reasoning steps. Our core idea is to integrate lightweight symbolic structures into few-shot prompts, making the inference steps more transparent and structured, while simultaneously strengthening the induction-head behavior of LLMs Olsson et al. (2022). Specifically, our Symbolic-Aided CoT prompting framework explicitly outlines essential reasoning substeps: rule matching—selecting constraint rules that align with the current state of inference, new premise inference—applying the selected rules to generate new premises, and knowledge base (KB) updating—appending the newly inferred premises to the KB. By incorporating these symbolic structures, our method preserves the flexibility and general applicability of standard prompting techniques while enhancing both the interpretability and analyzability of LLM reasoning behavior. Empirical evaluations across four reasoning QA benchmarks, ProofWriter, FOLIO, ProntoQA, and LogicalDeduction, demonstrate the effectiveness of our approach, particularly in scenarios involving complex reasoning that requires navigating multiple rules and constraints. Remarkably, our Symbolic-Aided CoT, when applied with open-source LLMs (e.g., Qwen3), achieves performance comparable to that of a complex multi-turn reasoning framework Sun et al. (2024) built on the powerful GPT-4 model, particularly on the ProofWriter, ProntoQA, and LogicalDeduction datasets.
@@ 2.1 Background
In the second approach, the original logical reasoning problem is decomposed into a sequence of subtasks. Each subtask is solved individually, and the process iterates over multiple turns until a specified stopping condition is satisfied, ultimately producing the final result: $\mathrm{Loop([\mathrm{LLM}^{\textrm{subtask}_{t}}(\mathcal{R},Q)]_{t})}$ where $\textrm{subtask}_{t}$ may represent an arbitrary LLM-based unit function such as rule matching, rule inference, or new premise scoring, among others. These sub‑tasks are typically carefully designed and arranged with sequential dependencies, with the aim of mitigating hallucinations.
@@ 2.2 Related Works
In this line of work, the main challenge lies in transforming the description of constraints into logical formulas, $\mathrm{LLM}^{\textrm{lang2logic}}$, effectively functioning as a semantic parser. More specifically, Yao et al. (2023) introduced a method to enrich CoT prompting by leveraging results from external APIs, invoking functions to retrieve supplementary information. Building on this idea, Gao et al. (2023) proposed an approach that augments the intermediate reasoning steps of CoT reasoning through a runtime environment (e.g., a Python interpreter), which has proven particularly effective for mathematical reasoning tasks. Further, Pan et al. (2023); Ye et al. (2023); Olausson et al. (2023) enhanced the performance of logical reasoning by translating all constraints in a sample into logical forms, completing a logical program, and solving it using an independent symbolic reasoner. In addition, Xu et al. (2024a) strengthened the logical‑form parsing process of the open-source LLMs by leveraging fine-tuning on a large‑scale, curated dataset. In contrast, our method does not rely on any external symbolic solver; instead, it integrates symbolic syntax directly into the reasoning process, aiming to enable the LLM itself to reason as a symbolic solver.
@@
This approach leverages LLMs directly by designing frameworks that decompose complex reasoning tasks into smaller, manageable subtasks—such as rule selection, premise derivation, and scoring—to enhance overall reasoning robustness (Zhang et al., 2025; Feng et al., 2024; Sun et al., 2024). In particular, Sun et al. (2024); Xu et al. (2024a) introduced a framework that enables LLMs to uncover hidden premises and integrates scoring components through a multi‑turn reasoning process. Similarly, Zhang et al. (2025) proposed a framework that incrementally refines reasoning via three subcomponents: Proposer, Verifier, and Reporter. Finally, Feng et al. (2024) presented LoGiPT, a method that enhances LLMs’ ability to function as logical solvers by learning the reasoning process step by step through additional training on large‑scale data collected from the reasoning traces of external symbolic solvers. Compared to our work, LoGiPT similarly performs step‑by‑step reasoning as a logical solver and can produce a proof tree at the end; however, our method achieves competitive performance without the need for any additional training.
@@ 3 Methodology
In this section, we present our Symbolic‑Aided CoT method, its variants, and the motivation behind it in comparison with baseline prompting techniques: Standard - which directly provides an answer without any reasoning - and CoT (Wei et al., 2022) - which produces an answer accompanied by step‑by‑step reasoning. All of these prompting methods are augmented with a hard-selected few‑shot examples included in the prompt Qi et al. (2025). All prompts are directly fed forward through the LLM to obtain the final predicted answer in a single turn:
$A^{\text{out}}=\mathrm{LLMs}(\mathrm{prompting}(\mathcal{R},Q))$ (1)
In this setup, the LLM is solely responsible for generating the desired answer given the contextual input. The prompting component consists of a few‑shot template designed to help the LLM understand the task description while leveraging its own knowledge to reason over the list of provided rules in the contextual information. For clarity, the templates for Standard and CoT prompting are shown in the first two rows of Table 1.
@@ 3.1 Symbolic-Aided CoT
We formulate logical reasoning tasks into three fundamental sub‑tasks, namely, reasoning operators: rule matching, rule inference, and knowledge base updating. Previous frameworks, such as DetermLR Sun et al. (2024) and CR Zhang et al. (2025), were also built on carefully designed unit operators, integrating them with procedural programming to process the outputs of LLMs. The key difference between our Symbolic‑Aided CoT and these approaches is that Symbolic‑Aided CoT is conceived entirely as an LLM‑driven program. In our design, the LLM is expected to learn the flow of the logical reasoning process from a few‑shot demonstration. To this end, the LLM has full visibility of all sub‑reasoning steps and autonomously decides which step to execute next. The overview of our Symbolic‑Aided CoT is presented in the third row of Table 1, which illustrates the instruction text, list of rules, question, and reasoning‑flow examples. For a clearer explanation, we elaborate on the two gray blocks shown in this table, which pertain to rule tagging and reasoning operators in our Symbolic‑Aided CoT.
@@
This demonstration serves as the core example that enables LLMs to learn, in context, the pattern for solving logical reasoning tasks. We use a set of symbols, similar to those in programming languages, to represent the inference flow (see Figure 2). At each reasoning step, the LLM selects the relevant rules and premises from the current knowledge base (KB) to infer new premises. Each newly inferred premise is then appended to the KB for use in the next inference step. A breadth‑first search strategy is applied to traverse the nodes (premises), as illustrated in Figure 2. Inspired by how humans solve such tasks, we maintain a KB state to prevent cyclical inference loops: if a newly inferred premise already exists in the KB, it is not added again. All patterns of rule selection, inference, and KB updating are implicitly conveyed within the demonstrations provided in the few‑shot prompts, allowing the LLM to internalize these reasoning steps.
Figure 2: Left: Graphical model of inference rules. Right: Reasoning flows in the Symbolic‑Aided CoT demonstrations.
@@ 4.1 Experimental Setup
We conducted our experiments primarily on open-source LLMs, including Llama-3.1-8B-Instruct Grattafiori et al. (2024) and the Qwen-3 models Yang et al. (2025). Specifically, we aim to evaluate the effectiveness of our Symbolic-Aided CoT compared to standard CoT and standard prompting on these LLMs. In addition, we performed extensive experiments on the Qwen-3 models across various sizes—1.7B, 4B, 8B, and 32B—to further assess the scalability and effectiveness of our proposed prompting method. Finally, for comparison with previous approaches, we also conducted extensive experiments on a powerful closed-source LLM, GPT-4 Achiam et al. (2023), to evaluate our proposed method. For open-source models, we use greedy decoding to generate the answer i.e., the token with the maximum logit is picked.
@@ 4.2 Main Results
These results demonstrate the clear superiority of our proposed method over the standard CoT on three datasets, ProofWriter, ProntoQA, and LogicalDeduction. Notably, on the ProofWriter dataset, Symbolic-Aided CoT significantly outperforms CoT, achieving improvements of $15\%$ and $1\%, 22$$2\%$ on the Qwen-3-8B, Qwen-3-14B, and Llama3.1-8B-Instruct models, respectively. In addition, the improvement is clearly observed on the LogicalDeduction and ProntoQA datasets across three open-source LLMs. These findings further highlight that the degree of improvement varies across different LLMs. The effectiveness of Symbolic-Aided CoT largely depends on each model’s ability to understand logical relationships and recognize logical matching patterns embedded within the few-shot prompts. Moreover, our method is simple yet effective, achieving competitive performance compared to prior works such as Logic-LM Pan et al. (2023) and DetermLR Sun et al. (2024), even when those methods are supported by a more powerful GPT-4 model.
@@
Finally, we evaluate our proposed method in the setting that uses GPT-4 as the backbone LLM for the reasoning task (last row of Table 2). Compared to the SymbCoT framework Xu et al. (2024a), our method achieves superior performance on ProntoQA but lower performance on other datasets. This difference can be attributed to SymbCoT’s use of complex interactive reasoning sub-steps—such as translator, planner, solver, and verifier—each supported by carefully designed prompts tailored to the specific sub-step and logical reasoning task. Overall, our method surpasses the performance of previous methods on the ProntoQA dataset and achieves remarkable results on the ProofWriter, FOLIO, and LogicalDeduction datasets. These results demonstrate the robustness of our approach, even in the stringent setting that uses only non-interactive inference without the support of an external solver or multi-step inferences, such as the DetermLR Sun et al. (2024) or LogicLM Pan et al. (2023) approaches.
@@
Fine-tuned ID Qi et al. (2025) Supervised FT Non-Interactive $71.67$ $70.00$
Standard ICL few-shot Non-Interactive $36.83$ $53.92$ $51.80$ $40.67$
CoT ICL few-shot Non-Interactive $44.83$ $\mathbf{56.86}$ $74.00$ $58.00$
Symbolic-Aided CoT (ours) ICL few-shot Non-Interactive $\mathbf{68.67}$ $55.88$ $\mathbf{89.00}$ $\mathbf{59.33}$
@@
Standard ICL few-shot Non-Interactive $60.00$ $62.25$ $80.80$ $57.00$
CoT ICL few-shot Non-Interactive $57.83$ $\mathbf{66.67}$ $95.80$ $72.67$
Symbolic-Aided CoT (ours) ICL few-shot Non-Interactive $\mathbf{78.67}$ $65.69$ $\mathbf{97.20}$ $\mathbf{77.33}$
@@
Standard ICL few-shot Non-Interactive $46.50$ $67.16$ $77.80$ $61.67$
CoT ICL few-shot Non-Interactive $62.67$ $\mathbf{74.02}$ $97.20$ $81.67$
Symbolic-Aided CoT (ours) ICL few-shot Non-Interactive $\mathbf{77.00}$ $65.20$ $\mathbf{97.80}$ $\mathbf{86.33}$
@@
CoT Sun et al. (2024) ICL few-shot Non-Interactive $67.41$ $67.65$ $91.00$ $73.33$
Logic-LM Pan et al. (2023) ICL few-shot Interactive +E.Solver $79.66$ $78.92$ $83.20$ $87.63$
DetermLR Sun et al. (2024) ICL few-shot Interactive +Programming $79.17$ $75.49$ $98.60$ $85.00$
SymbCoT Xu et al. (2024a) ICL few-shot Interactive +Programming $\mathbf{82.50}$ $\boldsymbol{83.33}$ $99.60$ $\mathbf{93.00}$
Symbolic-Aided CoT (ours) ICL few-shot Non-Interactive $77.09$ $74.51$ $\boldsymbol{100.00}$ $86.33$
@@
Furthermore, KB-tracking intuitively provides additional features to the hidden representation of premises, allowing the model to distinguish inferred premises from conditional premises in the constraint rules. In another aspect, the symbolic $\mathrm{Validate}$ function in our Symbolic-Aided CoT helps LLMs refer back to the original question in the context to select the appropriate premise for logical matching and producing the final answer.